Determine the t-value based on the sample size n and a

1)n=25 a= 0.01

Df=

T-Value =


2)n=15 a= 0.025

Df=

T-Value =


1
Expert's answer
2022-03-01T13:32:00-0500

1) "\alpha=0.01, df=n-1=24" degrees of freedom.


"t=2.228138"

2) "\alpha=0.025, df=n-1=14" degrees of freedom.


"t=2.509569"